golban:

There are 6 and 2 over 3pounds of walnuts in a container, which will be divided equally into containers that hold 1 and 1 over 3pounds. This would fill ____ and 9 over 12containers. Numerical Answers Expected!

golban:

help justic

golban:

hurry

golban:

pls

justus:

Okay

justus:

6 2/3 = 1 1/3 x

justus:

20/3 = 4/3 x

justus:

x = 20/3 * 3/4

justus:

x=5

justus:

This would fill __5__ and 9 over 12 containers.

justus:

Did you understand?

golban:

no

golban:

what is the awnser

justus:

I wrote it above

golban:

say it again pls

justus:

This would fill __5__ and 9 over 12 containers.

justus:

You got it?

justus:

You there?